OpenStudy (anonymous):

When a 20 V battery is hooked up to a 100 Ω resistor in a circuit, how much current flows through the circuit? A. 0.1 A B. 0.2 A C. 2.0 A D. 50 A

OpenStudy (michele_laino):

here, current is: I=V/R=20/100=...amperes

OpenStudy (anonymous):

0.2! so that is our solution?

OpenStudy (johnweldon1993):

^Correct

OpenStudy (anonymous):

ooh yay! thanks!!:D